Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-295_2000-301

Mission Operations Days: 2000/295 to 2000/301
October 21, 2000 00:00:00 GMT to October 27, 2000 00:00:00 GMT


Science Data Loss From SFE-A Anomaly (2000/300 14:05:02 to 2000/301 14:06:19) See Anomaly section below for details.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 5 sets of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 14H Gain Change (10 commands)
  • FR Band 24 Gain Change (10 commands)
  • FR Band 25 Gain Change (10 commands)
  • FR Band 27 Gain Change (10 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67, 68Nominal

This Week’s Completed MODIS Activities:

Saturday, October 21, 2000

None

Sunday, October 22, 2000

None

Monday, October 23, 2000

None

Tuesday, October 24, 2000

2000/298 22:05:22 Real-Time Cleared Inhibit ID #54 (SRCA Commands) (Set by MODIS SRCA TMON #66 trip on DOY 290)
2000/298 22:06:11 Real-Time Reset the SRCA Filter Wheel Step Count proc: MOD_SET_MTR_SW_COUNT(5,100)

The value of the SRCA Filter Wheel Step Count when the TMON #66 tripped was 100. The wheel was inhibited from being commanded (remaining at step 100), but the executing macro automatically sets the expected wheel position (Home, at step 20).

2000/298 23:03:43 Real-Time Returned SRCA Filter Wheel to Home (step 20) (Left at step 100 after the TMON #66 trip on DOY 290) proc: MOD_SRCA_HOME('N','RADIO')

Wednesday, October 25, 2000

None

Thursday, October 19, 2000

None

Friday, October 27, 2000

2000/301 09:14:13 ATC OA15: SDSM Open
2000/301 10:53:06 ATC OA16: SDSM Screened

The CAM for the MODIS Transition to B-Side Operations was approved (Friday, October 27, 2000). The transition is scheduled to begin at approximately 1300 GMT on Monday, October 30, 2000.

This Week's Scheduled MODIS Activities Not Completed:

Tuesday, October 24, 2000

2000/298 21:15:43 Real-Time Returned SRCA Filter Wheel to Home (step 20) proc: MOD_SRCA_HOME('Y','RADIO')

The parameters of this proc attempted to use a simple, software count method to return the SRCA motors to their home positions. The Filter Wheel's software count indicated it was currently at the Home position. Because of this, the SRCA Filter Wheel was not commanded to move. Before attempting this method again, the correct (non-Home) software step count of the Filter Wheel needed to be set.

2000/298 22:07:54 Real-Time Returned SRCA Filter Wheel to Home (step 20) proc: MOD_SRCA_HOME('Y','RADIO')

With the correct step count set, the Filter Wheel motor was commanded to move the necessary steps to return it to the home position (80 steps "backwards"). However, the Filter Wheel motor appears to have moved slightly slower than the software count was predicting. When the software count indicated that the Filter Wheel should be at the Home position, the wheel was still approximately 6 steps from Home (at step 26).

The MOD_SRCA_HOME proc was run for the 3rd time using the more robust method of moving the SRCA wheels (hence, the 'N' parameter). Instead of using software motor counts, this process instructs the SRCA Motors to continue turning until the Home triggers trip. At that point, the software step counts are updated to reflect the Home positions. This method is most useful when software counts have been lost/unknown, but also proved quite useful in this situation as well.

Upcoming MODIS Events:

Saturday, October 28, 2000

None

Sunday, October 29, 2000

None


Monday, October 30, 2000

Either:
2000/304 14:00:00 - 14:40:00 Real-time Begin Switch to B-side Science
2000/304 14:57:54 - 15:09:52 X-band service to Goddard
Or:
2000/304 15:21:18 - 15:59:58 Real-time Begin Switch to B-side Science (Note: MOPITT has requested commanding during this pass)
2000/304 16:35:01 - 16:48:47 X-band service to Goddard


The Switch to B-Side Science will nominally involve:

  • Turn off MODIS
  • Turn on MODIS with Power Supply 2 and Control Processor B (red limits for 2 minutes)
  • Disable TMON #68 commanding, FIFO double write
  • Transition to Science Mode on the B-side (red limits for 2 minutes) (may see TAXI errors for 15 minutes)
  • Enable TMON #68 commanding, FIFO double write
  • Verify Correct CP status words
  • Reset Formatter to Upload Mode
  • Load Formatter Patch #1
  • Load Formatter Patch #2
  • Restart Formatter with ASTATE = 0

If B-Side Transition occurs successfully, then perform the following initial calibrations:

Tuesday, October 31, 2000

2000/305 19:?? - 20:?? Real-Time Set SMIR Itwk/Vdet to 79/110
2000/305 20:21:52 ATC OA15: SDSM Open
2000/305 22:00:45 ATC OA16: SDSM Screened
2000/305 23:00:00 ATC Set Black Body Duty Cycle to 'FULL'
2000/305 23:00:30 ATC Set Black Body Temperature to '315K'

Wednesday, November 1, 2000

2000/306 06:00:00 ATC Set Black Body Duty Cycle to 'THIRD'
2000/306 06:00:30 ATC Set Black Body Temperature to '270K'

Thursday, November 2, 2000

2000/307 04:00:30 ATC Set Black Body Temperature to '290K'

Further calibration activities, gain changes, and Itwk/Vdet settings will be determined/proposed as the results of DOY 304 activities are finalized.

Maneuvers:

The next delta-v maneuver is scheduled for November 1st, 2000.

MODIS Anomalies:

On DOY 2000/300 (Thursday, October 26, 2000), the AM-1 Terra SFE-A experience an error causing TMON16 to trip, thereby shutting off SFE-A. Until the SFE-A turned back on, no Science Data (for any instrument) will be collected (full science data loss). See the Non-MODIS Significant Events section below for a full description of the event.

TMON #66, the MODIS SRCA A-side lamp current/voltage monitor tripped on day 290 shortly after beginning the 30W lamp portion of the macro. The TMON set the lamp usage to OFF, turned off the SRCA, and set inhibit ID #54 (MODIS SRCA commands). Because of this, the SRCA Filter wheel was left at step 100 rather than the home position.

A transient high current to the lamps appears to be a normal characteristic of the lamps. At this time, investigation is continuing to determine a possible correction to the TMON algorithm or alternate SRCA operating philosophy.

[DOY 301 update] Recent analysis indicates that the initial transient current spikes associated with an On-Orbit 30W test will consistently exceed the time out sampling (two or more readings exceeding the specified limit) for triggering the SRCA TMONs. Additionally, the MODIS Team at SBRS is confident that the over current/voltage protection built into the SRCA will provide greater protection than the current SRCA TMONs are capable of providing. This translates into the protection of the TMONs as only being a semi-effective "fall-back". The conclusion is for TMONs #66 and #67 to be disabled via an Engineering Work Order prior to the next SRCA activities.

Continued monitoring for formatter errors and loss of cooler margin.

Attached to this report is a spreadsheet listing the times, addresses, formatter temperature, satellite ground track, and space craft day/night position for a select number of the formatter events that have occurred since DOY 256. The current event count is 2576. Due to the high number of recent events, not all details have been updated in the anomaly log.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches.

The upcoming week will see the attempted switch to B-Side Operations for MODIS. A few of the major goals this switch will hope to produce:

  • A more complete determination as to the cause of the sporadic A-Side Formatter Error Events.
  • MCST's preferred Science Configuration for the upcoming "Golden Month" of November Science Data Collection.

MODIS Telemetry Trends:

The Radiative Cooler intermediate stage temperature has slowly been increasing since approximately day 240. During this time the LWIR focal plane heater margin has been decreasing. The IOT is continuing to trend these telemetry mnemonics.

Non-MODIS Significant Events:

The following message was distributed in regards to the SFE-A Anomaly:

"...the SFE-A shut off at 300/14:05:02. The anomaly investigation has been on going and the team has determined that the HPM2 card failed to send telemetry to the LPM. As a result TMON16 detected the HPM2BIT status as failed and sent the command to turn off SFE-A. No science data will be available until the SFE functionality is restored."

The anomaly resolution team concluded that the error was probably caused by a single event upset. The SFE-A was restarted, and science data collection appears to be proceeding nominally.

The duration that the SFE-A was disabled: 2000/300 14:05:02 to 2000/301 14:06:19 (Just over 24 hours).

NO Science Data will be available for this time period. All Science Data in this time frame has been permanently lost.

Limited Life Item Status:

SRCA 10W Lamp #1: 173.5 of 500 hours
SRCA 10W Lamp #2: 135.9 of 500 hours
SRCA 10W Lamp #3: 145.1 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 556.1 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1447 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements